the sum of two numbers is 10. twice the larger number minus the smaller number is 15 . what are the numbers?
Answer by ewatrrr(10682)


Hi
sum of two numbers is 10
Let x and (10-x) represent the larger and smaller number respectively
Question States***Write as we Read
2x - (10-x) = 15
Solving for x
2x - 10 + x = 15
3x = 25
x = 25/3 or 8 1/3, the larger. The smaller 1 2/3 or 5/3 (10- 8 1/3)
CHECKING our Answer
2*(25/3) - 5/3 = 50/3 - 5/3 = 45/3 = 15